OpenStudy (anonymous):

a person earned $20 on 1 day and spent $ 15 on next day. He again earned $20 on the 3rd day and spent $15 on the fourth day and so on. On which day he will first have $60 with him?

OpenStudy (lgbasallote):

first im going to ask you..if he earns 20 one day and loses 15 the next day...how much does he have in 2 days?

OpenStudy (anonymous):

5+20+20=25?

OpenStudy (anonymous):

Every 2 day that person would have $5. just keep adding 'till you get $60

OpenStudy (inkyvoyd):

There's a major problem with this problem. It doesn't say how much money he starts out with.

OpenStudy (inkyvoyd):

@znUnz , no, actually

OpenStudy (lgbasallote):

huh? that's 20 - 15 in two days

OpenStudy (inkyvoyd):

at 40 dollars, he will get to 60.

OpenStudy (anonymous):

k sry i thoght that he s earning 20 daily and not spending any

OpenStudy (lgbasallote):

so how much does he have in 2 days?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

on day 24 you will get $60

OpenStudy (lgbasallote):

60 divided by 5?

OpenStudy (anonymous):

20-15=5 therefore he takes two days to spent...therefore 5x=60*2 x=60*2/5 x=24 24 days

OpenStudy (anonymous):

the ans is 17th day

OpenStudy (inkyvoyd):

@aroo1993

OpenStudy (anonymous):

may be the ans given in my book is wrong........

OpenStudy (inkyvoyd):

On day 16 you have 40, on day 17 you have 60. I've been trying to tell you guys that the whole time.

OpenStudy (anonymous):

17 day seems correct

OpenStudy (anonymous):

Yes 16 to 17 is correct as inky stated

OpenStudy (inkyvoyd):

I tried to explain that to some of you, but I'm not sure I did it very well :/

OpenStudy (anonymous):

( I got it ;-)

OpenStudy (inkyvoyd):

:)

OpenStudy (anonymous):

how cum?

OpenStudy (lgbasallote):

^LOL

OpenStudy (inkyvoyd):

Well, how much money do you have on day 16, @aroo1993

OpenStudy (anonymous):

40

OpenStudy (inkyvoyd):

And how much money do you have on day 17?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

You get paid on 17

OpenStudy (anonymous):

k thnx

OpenStudy (inkyvoyd):

Let me explain. Say a frog jumps 99 feet and drops 98 feet right after. If a wall is 100 feet, how many jumps does it take him to go as high as the wall?

OpenStudy (inkyvoyd):

Only 2.

OpenStudy (inkyvoyd):

Not 100, because on the second jump, he is already 100 feet, even though he falls back to 2 feet.

OpenStudy (anonymous):

guys its wrong i think i may hav the answer 20-15=5 therefore he takes one and half days to spent...therefore 5x=60*2 x=60*1.5/5 x=18 18 days

OpenStudy (lgbasallote):

do it manually day 1 - 20 2 - 5 3 - 25 4 - 10 5 - 30 6 - 15 7 - 35 8 - 20 9 - 40 10 - 25 11 - 45 12 - 30 13 - 50 14 - 35 15 - 55 16 - 40 17 - 60 see?

OpenStudy (inkyvoyd):

18-45 19-65 20-50 *THESE NUMBERS DO NOT MATTER BECAUSE WE ALREADY REACHED 60 DOLLARS ON DAY 17!*

OpenStudy (anonymous):

Wait. If you're talking about money saved then by the 24th day he would have saved $60. If you want to have $60, then you would have $60 by the 17th day. ON the 18th day however, it drops back down to $45 Is that right @inkyvoyd

OpenStudy (inkyvoyd):

Exactly, @znUnz